# Universal Gröbner Bases of

### v2

@@ 2.1. Simplicial Complexes and Matroids
Next we recall some basic aspects of Stanley-Reisner theory —see eg. [8, 11, Ch. 1] for more details. The Stanley-Reisner ideal of a simplicial complex $\Delta$ on $S$ is the monomial ideal $J_{\Delta}$ in ${\mathbb{C}}[x_{i}:i\in S]$ generated by the square-free monomials $x^{A}:=\Pi_{i\in A}x_{i}$ such that $A$ is a nonface of $\Delta$, i.e., $A\not\in\Delta$. If $\Delta$ is pure then $J_{\Delta}$ is equidimensional, i.e., all its associated prime ideals have the same dimension. On the other hand, every ideal $J$ in ${\mathbb{C}}[x_{i}\,:\,i\in S]$ generated by square-free monomials is the Stanley-Reisner ideal of a simplicial complex $\Delta_{J}$ on $S$; $x^{A}\in J$ if and only if $A$ is a nonface of $\Delta_{J}$. The inclusion-minimal nonfaces of $\Delta_{J}$ are in bijection with the minimal generators of $J$. The simplicial complex $\Delta_{J}$ is called the Stanley-Reisner complex of $J$.
@@
[9, Theorem 12.3.1] The union of matroids $\mathcal{M}=\mathcal{M}_{1}\vee\cdots\vee\mathcal{M}_{n}$ is a matroid, whose rank function $r_{\mathcal{M}}$ can be related to the rank functions $r_{1},\ldots,r_{n}$ of $\mathcal{M}_{1},\ldots,\mathcal{M}_{n}$ as follows:
@@ 2.2. Universality Criterion for Gröbner Bases
There is a natural embedding of $V$ in the projective space $({\mathbb{P}}^{1})^{N}$ obtained by homogenizing each coordinate $x_{i}$, i.e., sending $x_{i}\mapsto[x_{i}:1]$. Let $\overline{V}$} denote the closure of $V$ inside $({\mathbb{P}}^{1})^{N}.$ Given a subset $U\subset[N],$ consider the projection
@@
To apply Theorem 2.5, we must show that the projection ${\overline{\pi}}_{U}:{\overline{V}}\to({\mathbb{P}}^{1})^{U}$ is for each facet of the simplicial complex ${\Delta}(f_{1},\ldots,f_{r}).$ Since our varieties are affine, it will be convenient to instead consider the projections $\pi_{U}:V\to({\mathbb{C}}^{1})^{U}.$ The following simple lemma expresses the the surjectivity, and hence dominance, of ${\overline{\pi}}_{U}$ in terms of the original affine variety $V$, when $V$ is irreducible.
@@
In order to show that $\{f_{1},\ldots,f_{r}\}$ is a Gröbner basis of $\mathcal{I}(V)$ with respect to any monomial order on ${\mathbb{C}}[x_{1},\ldots,x_{N}]$, it suffices to argue that $\{f_{1}^{h},\ldots,f_{r}^{h}\}$ is a Gröbner basis for $\mathcal{I}(\overline{V})$ with respect to any product order such that $x_{j+N}<x_{i}$ for all $i,j.$ These orders can be viewed as natural extensions of monomial orders on ${\mathbb{C}}[x_{1},\ldots,x_{N}]$ to monomial orders on ${\mathbb{C}}[x_{1},\ldots,x_{2N}].$ If $f_{1}^{h},\ldots,f_{r}^{h}$ form a Gröbner basis of ${\mathcal{I}}({\overline{V}})$ with respect to <,<, then $f_{1},\ldots,f_{r}$ form a Gröbner basis of ${\mathcal{I}}(V)$ with respect to the corresponding monomial order on ${\mathbb{C}}[x_{1},\ldots,x_{N}],$ see for example, [4, Exercise 15.39].
@@
Since the variety $V$ is irreducible, the sets $U\subset[N]$ with ${\overline{\pi}}_{U}({\overline{V}})=({\mathbb{P}}^{1})^{U}$ are precisely the independent sets in an algebraic matroid [9, §6.7], whose rank is $\operatorname{dim}V.$ The bases in this matroid correspond to transcendence bases of rational functions in the set $\{x_{1},\ldots,x_{N}\}\subset{\mathbb{C}}(V)$. As in [10], we call this the algebraic matroid of $V.$
@@
Returning to the proof of Theorem 2.5, we now prove that condition (2) holds, i.e., that $J=\langle\operatorname{in}_{<}(f_{1}^{h}),\ldots,\operatorname{in}_{<}(f_{r}^{h})\rangle$ is equidimensional, using Stanley-Reisner theory. Observe that the Stanley-Reisner complex $\Delta_{J}$ is closely related to the simplicial complex $\Delta(f_{1},\ldots,f_{r})$; for any $f\in{\mathbb{C}}[x_{1},\ldots,x_{N}]$ there is a bijective correspondence between ${\rm spr}f$ in $[N]$ and the support of $\operatorname{in}_{<}(f^{h})$ in $[2N]$. Hence, the nonfaces of ${\Delta}_{J}$ and $\Delta(f_{1},\ldots,f_{r})$ are in bijection. Moreover, by [6, Lemma 2.10], there is a bijection sending facets $U\in\Delta(f_{1},\ldots,f_{r})$ of size $k$ to facets $\widetilde{U}\in\Delta_{J}$ of size $k+N.$
@@
([6, Lemma 2.10]) $=N+\operatorname{dim}\Delta(f_{1},\ldots,f_{r})+1$
@@
(See also [7, Proposition 1.7.3], for a more general statement.) Thus, we have
@@ 4.1. The simplicial complex
Let $U$ be any facet of $\Delta_{n}$. Whether $U$ has profile $(3,2,1,1,\ldots,1)$ or $(2,2,2,1,\ldots,1)$, it must must contain at least one element from $\{x_{i1},x_{i2},x_{i3}\}$ for each $i=1,\ldots,n$. In other words, $U$ must contain a basis $B$ of the direct sum $\mathcal{U}_{1,n}^{\oplus n}$. Now, since the remaining 3 elements of $U\setminus B$ form a basis of $\mathcal{U}_{3,3n}$, we have that $U$ is a basis in $\mathcal{U}_{3,3n}\vee\mathcal{U}_{1,n}^{\oplus n}.$ A similar proof shows that any face of $\Delta_{n}$ is a union of independent sets in $\mathcal{U}_{3,3n}$ and $\mathcal{U}_{1,n}^{\oplus n}.$ Since the class of transversal matroids is closed under union ([9, Corollary 12.3.8]), this completes the proof. ∎
@@
In matroid theory, it is well-known that every transversal matroid is representable over a sufficiently large (finite) field (cf. [9, Ch. 6].) On the other hand, the matroids $\Delta_{n}$ are not realizable over arbitrary fields. For example, starting from $\Delta_{4}$, we may delete the variables $x_{21},x_{31},x_{41}$ and contract the variables $x_{11},x_{12},x_{22},x_{32},x_{42}$ to obtain the uniform matroid $\mathcal{U}_{2,4}$ on the remaining variables $x_{13},\ldots,x_{43}$ as a minor. This implies that $\Delta_{n}$ for $n\geq 4$ is not representable over the field with two elements. In particular, the matroid $\Delta_{n}$ is non-graphic [9, Chapter 5].
@@ 5. The universal multiview ideal
For $n$ unknown cameras, the universal multiview variety consists of all points $(A_{1},\ldots,A_{n},p_{1},\ldots,p_{n})\in({\mathbb{P}}^{3\times 4})^{n}\times({\mathbb{P}}^{2})^{n}$ such that for each $i=1,\ldots,n$ there is some world point $q\in{\mathbb{P}}^{3}$ such that $A_{i}q=p_{i}.$ Let ${\widetilde{M}}_{n}$ be the affine cone over the universal multiview variety and ${\widetilde{I}}_{n}$ the vanishing ideal of ${\widetilde{M}}_{n}.$ Note that the difference between $M_{n}$ and ${\widetilde{M}}_{n}$ is that the former lies in $({\mathbb{P}}^{2})^{n}$ and its vanishing ideal $I_{n}$ lies in ${\mathbb{C}}[x_{1},\ldots,x_{n}]$ where $x_{i}$ consists of the three variables corresponding to the $i$th camera plane, while ${\widetilde{M}}_{n}$ lies in $({\mathbb{P}}^{3\times 4})^{n}\times({\mathbb{P}}^{2})^{n}$ and its vanishing ideal ${\widetilde{I}}_{n}$ lies in the polynomial ring with $12n$ variables coming from the $n$ cameras $A_{1},\ldots,A_{n}$ and the $3n$ variables of the image plane coordinates $x_{1},\ldots,x_{n}$.
@@ 5.1. The simplicial complex
with $U^{\prime}$ and $U^{\prime\prime}$ as in the preceding paragraph, and similarly for $\ell\geq 2.$ Finally, our observations preceding the proof of Theorem 5.2 imply that all facets of the matroid union $\mathcal{U}_{3,15n}\vee\mathcal{U}_{13,15}^{\oplus n}$ have the form $U_{\ell}$ for some choice of $U_{0}$ and $\ell.$
@@ 5.2. Applying the Huang-Larson theorem
Choose a generic point $Q\in({\mathbb{C}}^{1})^{U}$. We will argue that there is a point $P\in{\widetilde{M}}_{n}$ such that $\pi_{U}(P)=Q$. We consider a generic point $Q^{\prime}\in({\mathbb{C}}^{1})^{U^{\prime}}$ obtained from $Q$ by deleting the entry in position $x_{ij}$ and adding an entry in position $a_{ijk}$. By assumption, there is a point $P^{\prime}\in{\widetilde{M}}_{n}$ such that $\pi_{U^{\prime}}(P^{\prime})=Q^{\prime}$. The projection $\pi_{U}(P^{\prime})$ agrees with $Q$ in all positions except the one indexed by $x_{ij}$. Note that the entry of $P^{\prime}$ in position $a_{ijk}$ is irrelevant for the projection $\pi_{U}$ since $U$ does not contain $a_{ijk}$. We will use this flexibility to construct a $P\in{\widetilde{M}}_{n}$, from $P^{\prime}$, such that $\pi_{U}(P)=Q$.
@@ 6. Conclusion
This paper demonstrates that Theorem 2.5 is a useful tool for proving universal Gröbner basis results. We are optimistic that the techniques developed here can be adapted to other problems in the emerging field of algebraic vision, in particular, to several other families of ideals that appear in [1].

### v3

@@ 2.1. Simplicial Complexes and Matroids
Next we recall some basic aspects of Stanley-Reisner theory —see eg. [10, 13, Ch. 1] for more details. The Stanley-Reisner ideal of a simplicial complex $\Delta$ on $S$ is the monomial ideal $J_{\Delta}$ in ${\mathbb{C}}[x_{i}:i\in S]$ generated by the square-free monomials $x^{A}:=\Pi_{i\in A}x_{i}$ such that $A$ is a nonface of $\Delta$, i.e., $A\not\in\Delta$. If $\Delta$ is pure then $J_{\Delta}$ is equidimensional, i.e., all its associated prime ideals have the same dimension. On the other hand, every ideal $J$ in ${\mathbb{C}}[x_{i}\,:\,i\in S]$ generated by square-free monomials is the Stanley-Reisner ideal of a simplicial complex $\Delta_{J}$ on $S$; $x^{A}\in J$ if and only if $A$ is a nonface of $\Delta_{J}$. The inclusion-minimal nonfaces of $\Delta_{J}$ are in bijection with the minimal generators of $J$. The simplicial complex $\Delta_{J}$ is called the Stanley-Reisner complex of $J$.
@@
[11, Theorem 12.3.1] The union of matroids $\mathcal{M}=\mathcal{M}_{1}\vee\cdots\vee\mathcal{M}_{n}$ is a matroid, whose rank function $r_{\mathcal{M}}$ can be related to the rank functions $r_{1},\ldots,r_{n}$ of $\mathcal{M}_{1},\ldots,\mathcal{M}_{n}$ as follows:
@@ 2.2. Universality Criterion for Gröbner Bases
There is a natural embedding of $V$ in the multiprojective space $({\mathbb{P}}^{1})^{N}$ obtained by homogenizing each coordinate $x_{i}$, i.e., sending $x_{i}\mapsto[x_{i}:1]$. Let $\overline{V}$} denote the closure of $V$ inside $({\mathbb{P}}^{1})^{N}.$ Given a subset $U\subset[N],$ consider the projection
@@
To apply Theorem 2.5, we must show that the projection ${\overline{\pi}}_{U}:{\overline{V}}\to({\mathbb{P}}^{1})^{U}$ is surjective for each facet of the simplicial complex ${\Delta}(f_{1},\ldots,f_{r}).$ Since our varieties are affine, it will be convenient to instead consider the projections $\pi_{U}:V\to({\mathbb{C}}^{1})^{U}.$ The following simple lemma expresses the the surjectivity, and hence dominance, of ${\overline{\pi}}_{U}$ in terms of the original affine variety $V$, when $V$ is irreducible.
@@
In order to show that $\{f_{1},\ldots,f_{r}\}$ is a Gröbner basis of $\mathcal{I}(V)$ with respect to any monomial order on ${\mathbb{C}}[x_{1},\ldots,x_{N}]$, it suffices to argue that $\{f_{1}^{h},\ldots,f_{r}^{h}\}$ is a Gröbner basis for $\mathcal{I}(\overline{V})$ with respect to any product order such that $x_{j+N}<x_{i}$ for all $i,j.$ These orders can be viewed as natural extensions of monomial orders on ${\mathbb{C}}[x_{1},\ldots,x_{N}]$ to monomial orders on ${\mathbb{C}}[x_{1},\ldots,x_{2N}].$ If $f_{1}^{h},\ldots,f_{r}^{h}$ form a Gröbner basis of ${\mathcal{I}}({\overline{V}})$ with respect to <,<, then $f_{1},\ldots,f_{r}$ form a Gröbner basis of ${\mathcal{I}}(V)$ with respect to the corresponding monomial order on ${\mathbb{C}}[x_{1},\ldots,x_{N}],$ see for example, [5, Exercise 15.39].
@@
Since the variety $V$ is irreducible, the sets $U\subset[N]$ with ${\overline{\pi}}_{U}({\overline{V}})=({\mathbb{P}}^{1})^{U}$ are precisely the independent sets in an algebraic matroid [11, §6.7], whose rank is $\operatorname{dim}V.$ The bases in this matroid correspond to transcendence bases of rational functions in the set $\{x_{1},\ldots,x_{N}\}\subset{\mathbb{C}}(V)$. As in [12], we call this the algebraic matroid of $V.$
@@
Returning to the proof of Theorem 2.5, we now prove that condition (2) holds, i.e., that $J=\langle\operatorname{in}_{<}(f_{1}^{h}),\ldots,\operatorname{in}_{<}(f_{r}^{h})\rangle$ is equidimensional, using Stanley-Reisner theory. Observe that the Stanley-Reisner complex $\Delta_{J}$ is closely related to the simplicial complex $\Delta(f_{1},\ldots,f_{r})$; for any $f\in{\mathbb{C}}[x_{1},\ldots,x_{N}]$ there is a bijective correspondence between ${\rm spr}f$ in $[N]$ and the support of $\operatorname{in}_{<}(f^{h})$ in $[2N]$. Hence, the nonfaces of ${\Delta}_{J}$ and $\Delta(f_{1},\ldots,f_{r})$ are in bijection. Moreover, by [7, Lemma 2.10], there is a bijection sending facets $U\in\Delta(f_{1},\ldots,f_{r})$ of size $k$ to facets $\widetilde{U}\in\Delta_{J}$ of size $k+N.$
@@
([7, Lemma 2.10]) $=N+\operatorname{dim}\Delta(f_{1},\ldots,f_{r})+1$
@@
(See also [9, Proposition 1.7.3], for a more general statement.) Thus, we have
@@ 4.1. The simplicial complex
Let $U$ be any facet of $\Delta_{n}$. Whether $U$ has profile $(3,2,1,1,\ldots,1)$ or $(2,2,2,1,\ldots,1)$, it must must contain at least one element from $\{x_{i1},x_{i2},x_{i3}\}$ for each $i=1,\ldots,n$. In other words, $U$ must contain a basis $B$ of the direct sum $\mathcal{U}_{1,n}^{\oplus n}$. Now, since the remaining 3 elements of $U\setminus B$ form a basis of $\mathcal{U}_{3,3n}$, we have that $U$ is a basis in $\mathcal{U}_{3,3n}\vee\mathcal{U}_{1,n}^{\oplus n}.$ A similar proof shows that any face of $\Delta_{n}$ is a union of independent sets in $\mathcal{U}_{3,3n}$ and $\mathcal{U}_{1,n}^{\oplus n}.$ Since the class of transversal matroids is closed under union ([11, Corollary 12.3.8]), this completes the proof. ∎
@@
In matroid theory, it is well-known that every transversal matroid is representable over a sufficiently large (finite) field (cf. [11, Ch. 6].) On the other hand, the matroids $\Delta_{n}$ are not realizable over arbitrary fields. For example, starting from $\Delta_{4}$, we may delete the variables $x_{21},x_{31},x_{41}$ and contract the variables $x_{11},x_{12},x_{22},x_{32},x_{42}$ to obtain the uniform matroid $\mathcal{U}_{2,4}$ on the remaining variables $x_{13},\ldots,x_{43}$ as a minor. This implies that $\Delta_{n}$ for $n\geq 4$ is not representable over the field with two elements. In particular, the matroid $\Delta_{n}$ is non-graphic [11, Chapter 5].
@@ 5. The universal multiview ideal
For $n$ unknown cameras, the universal multiview variety consists of all points $(A_{1},\ldots,A_{n},p_{1},\ldots,p_{n})\in({\mathbb{P}}^{3\times 4-1})^{n}\times({\mathbb{P}}^{2})^{n}$ such that for each $i=1,\ldots,n$ there is some world point $q\in{\mathbb{P}}^{3}$ such that $A_{i}q=p_{i}.$ Let ${\widetilde{M}}_{n}$ be the affine cone over the universal multiview variety and ${\widetilde{I}}_{n}$ the vanishing ideal of ${\widetilde{M}}_{n}.$ Note that the difference between $M_{n}$ and ${\widetilde{M}}_{n}$ is that the former lies in $({\mathbb{P}}^{2})^{n}$ and its vanishing ideal $I_{n}$ lies in ${\mathbb{C}}[x_{1},\ldots,x_{n}]$ where $x_{i}$ consists of the three variables corresponding to the $i$th camera plane, while ${\widetilde{M}}_{n}$ lies in $({\mathbb{P}}^{3\times 4-1})^{n}\times({\mathbb{P}}^{2})^{n}$ and its vanishing ideal ${\widetilde{I}}_{n}$ lies in the polynomial ring with $12n$ variables coming from the $n$ cameras $A_{1},\ldots,A_{n}$ and the $3n$ variables of the image plane coordinates $x_{1},\ldots,x_{n}$.
@@ 5.1. The simplicial complex
with $U^{\prime}$ and $U^{\prime\prime}$ as in the preceding paragraph, and similarly for $\ell\geq 2.$ Finally, combining observations from Lemma 5.1 and Theorem 5.2, all facets of the matroid union $\mathcal{U}_{3,15n}\vee\mathcal{U}_{13,15}^{\oplus n}$ have the form $U_{\ell}$ for some choice of $U_{0}$ and $\ell.$
@@ 5.2. Applying the Huang-Larson theorem
Choose a generic point $Q\in({\mathbb{C}}^{1})^{U}$. We will argue that there is a point $P\in{\widetilde{M}}_{n}$ such that $\pi_{U}(P)=Q$. We consider a generic point $Q^{\prime}\in({\mathbb{C}}^{1})^{U^{\prime}}$ obtained from $Q$ by deleting the entry in position $x_{ij}$ and adding an entry in position $a_{ijk}$. By assumption, there is a point $P^{\prime}\in{\widetilde{M}}_{n}$ such that $\pi_{U^{\prime}}(P^{\prime})=Q^{\prime}$. The projection $\pi_{U}(P^{\prime})$ agrees with $Q$ in all positions except the one indexed by $x_{ij}$. Note that the entry of $P^{\prime}$ in position $a_{ijk}$ is irrelevant for the projection $\pi_{U}$ since $U$ does not contain $a_{ijk}$. We use this flexibility to construct $P\in{\widetilde{M}}_{n}$, from $P^{\prime}$ with $\pi_{U}(P)=Q$.
@@ 6. Conclusion
In conclusion, this paper demonstrates that Theorem 2.5 is a useful tool for proving universal Gröbner basis results in the emerging field of algebraic vision [8]. Specifically, we have obtained a new proof of the previously-known result Theorem 1.1, and an entirely new result in Theorem 1.2. We are optimistic that the techniques developed here can be adapted to other problems; in particular, to families of ideals appearing in [1], such as the resectioning varieties studied in [3].